# Magneto-optical properties of Group-IV–vacancy centers in diamond upon hydrostatic pressure

### v1

@@ III.4 Hyperfine interaction
We computed the hyperfine tensors for the relevant isotopes of dopants and proximate 13C $I=1/2$ nuclear spins that can be basically resonantly addressed optically Becker et al. (2016); Green et al. (2017). The 13C nuclear spins occur with 1.1% abundance in natural diamonds. The G4V(--) color centers are often generated by ion implantation for which the isotope can be well controlled. The non-zero nuclear spin isotopes are 29Si, 73Ge, 117Sn and 207Pb with $I=1/2$, $I=9/2$, $I=1/2$, and $I=1/2$, respectively. For 73Ge with $I=9/2$ the quadrupole moment also contributes to the fine level structure of the GeV(--) color center that we will also consider in this section. If these nuclear spins can be well controlled, then they could be applied as quantum memory of the system Metsch et al. (2019); Knaut et al. (2024) but may also contribute to extra spin flip-flop processes by reducing the coherence time of the electron spin of G4V(--) centers. Therefore, understanding the full hyperfine tensor and its effects on the system is crucial in quantum technology applications. As the electronic states are dynamically distorted by phonons, such hyperfine and quadrupole matrix elements are activated that play an important role in the possible spin relaxation processes Monge et al. (2023); Thiering and Gali (2024). In this study, we consider the full dynamical hyperfine and quadrupole tensors in the analysis of G4V(--) color centers.
We start the general description of the hyperfine interaction with taking into account the ${}^{2}E_{g,u}$ double degenerate electronic states, which leads to dynamical terms due to strong electron-phonon coupling,
$\hat{W}=\overleftarrow{S}\overleftrightarrow{A}\overrightarrow{I}+\overleftarrow{S}\overleftrightarrow{A}_{x}\overrightarrow{I}\hat{\sigma}_{z}+\overleftarrow{S}\overleftrightarrow{A}_{y}\overrightarrow{I}\hat{\sigma}_{x}$ (3)
where the $A$ is the well-known static term and $A_{x}$ and $A_{y}$ are the dynamic terms of the hyperfine tensor (see Appendix B and Refs. Monge et al., 2023; Thiering and Gali, 2024). The dopant ion in G4V(--) center sits in the inversion point of the defect, therefore, the electron spin and nuclear spin of the system can be simultaneously diagonalized, in other words, the double group symmetry of the static hyperfine interaction remains within $D_{3d}$ symmetry. In this case, the diagonal form of the static hyperfine tensor can be simplified to
$\hat{H}_{\text{HF-stat}}=\frac{1}{2}A_{\bot}\left(\hat{S}_{+}\hat{I}_{-}+\hat{S}_{-}\hat{I}_{+}\right)+A_{\|}\hat{S}_{z}\hat{I}_{z}\text{,}$ (4)
where $A_{\|}$ is parallel to the symmetry axis whereas $A_{\bot}$ term is able to induce spin flip-flop process between the electron spin - nuclear spin states within the same orbital component. However, the electron-phonon induced hyperfine interaction should also be considered that reads as
$\hat{H}_{\text{HF-dyn}}=$ $A_{1}\left[(\hat{S}_{+}\hat{I}_{z}+\hat{S}_{z}\hat{I}_{+})\hat{\sigma}_{-}+(\hat{S}_{z}\hat{I}_{-}+\hat{S}_{-}\hat{I}_{z})\hat{\sigma}_{+}\right]+$ (5)
$A_{2}\left[(\hat{S}_{-}\hat{I}_{-}\hat{\sigma}_{-}+\hat{S}_{+}\hat{I}_{+})\hat{\sigma}_{+}\right]\text{,}$
@@
where $A_{1}$ and $A_{2}$ parameters are defined in the Appendix B. It is important to highlight that $A_{1}$ and $A_{2}$ parameters are subject of the JT effect so the electronic parameters as computed within Born–Oppenheimer approximation will be reduced as explained in the Appendix B. We note that the orbital angular momentum is significantly quenched but cannot be fully neglected Thiering and Gali (2018). As a consequence, the remaining orbital angular momentum may contribute to the effective dipole-dipole hyperfine coupling. On the other hand, we assume that this contribution is small (e.g., few percents) and we do not consider it further in this context.
The strong electron-phonon coupling is responsible for relatively short coherence times of the electron spin in G4V(--) defects even at cryogenic temperatures Kepesidis et al. (2016), thus no electron spin resonance measurements have been reported for G4V(--) defects. The hyperfine structure may be observed by optical means. In the optical transition, the nuclear spin state does not flip thus it could be observed as an energy difference between the same $m_{I}$ hyperfine levels at ${}^{2}E_{u}$ ($A_{u}$) and ${}^{2}E_{g}$ ($A_{g}$) which is labeled as $A_{\text{PLE}}^{\text{exp}}=1/2(A_{u}-A_{g})$. In the case of $I=1/2$, two lines appear in the spectrum, which is then separated by $A_{\text{PLE}}^{\text{exp}}$. Typically, the chosen final state in the experiments is the lowest energy branch of ${}^{2}E_{3/2,g}$ because it has much longer coherence times than that of the upper branch. In many experiments, they resonantly excite the lower branch at ${}^{2}E_{3/2,u}$, so ${}^{2}E_{3/2,u}\rightarrow^{2}E_{3/2,g}$ optical transition occurs in the ZPL fluorescence. They typically apply relatively small magnetic fields (1.5–40 G) to split the Kramers-doublets that act as qubits: $E_{+,\uparrow,g}$ and $E_{-,\downarrow,g}$ of which levels are separated by $g_{g}\mu_{B}B$ where $g_{g}\approx 2$, $\mu_{B}$ and $B$ are the effective g-constant, the Bohr-magneton of the electron and the strength of the applied magnetic field, and the arrows represent the electron spin projections. The nuclear Zeeman shifts can be neglected beside these terms. Within the ${}^{2}E_{3/2,u}$ states the effect is similar that goes with $g_{u}\approx 2$ effective g-constant upon external magnetic fields. At the same time, similar splitting occurs in the upper branches (${}^{2}E_{-,\uparrow,g/u}$ and ${}^{2}E_{+,\downarrow,g/u}$) (see Fig. 4).
@@
Figure 4: The electron spin density of SiV(--) center with low $C_{2h}$ symmetry in the electronic (a) ground and (b) excited states is depicted with isosurface value of $5\times 10^{-3}$ $e/$Å. (c) Scheme of the hyperfine splitting and EPR spectrum for a system with electron and nuclear spin of 1/2. (d) The possible electronic transition between two excited and ground state within the hyperfine splitting are depicted in blue and orange lines. Hence, the optical spacing $A_{\text{PLE}}$ can be detected.
@@
Table 3: The hyperfine parameters of the G4V(--) centers as obtained by HSE06 functional in $C_{2h}$ symmetry configuration. The ($A_{\|}$, $A_{\bot}$) and ($A_{1}$, $A_{2}$) parameters are defined in Eqs. (4) and (5), respectively, where we list here the reduced parameters with reduction factors given in Table 7. The calculated $A_{\text{PLE}}$ values are valid at zero magnetic field as observed in the optical transition between the respective $E_{3/2,u}$ and $E_{3/2,g}$ states (see text for details). $A_{\text{PLE}}^{\text{exp}}$ data are taken with different experiments at various external magnetic fields. Our theory implies that the applied magnetic field in the range of 10-50 Gauss has less than 0.1% effect on $A_{\text{PLE}}$).
@@
29Si 1/2 0.03 $-0.01$ 4.8 5.1 - -
73Ge 9/2 $-0.22$ 0.26 4.5 4.6 - -
117Sn 1/2 $-0.1$ $-0.2$ $29.9$ $32.3$
207Pb 1/2 $-0.06$ 0.4 $-19.4$ $-20.76$
In Table 3 we list the calculated hyperfine parameters of dopants for G4V(--) centers. In these calculations, we use HSE06 functional in order to obtain accurate results for the spin density near the nuclear spins. We find that the spin density cloud around the dopant ion is less dense in the electronic excited state than that in the electronic ground state in each G4V(--) color center. As a consequence, the static hyperfine parameters of dopant ion are larger in the electronic ground state than that in the electronic excited state. As a consequence, the hyperfine lines for $I>0$ dopants can be indeed observed because $A_{\text{PLE}}$ becomes non-zero.
It can be also seen that the dynamic hyperfine parameters are small but non-zero in both ${}^{2}E_{u}$ and ${}^{2}E_{g}$ states as listed in Table 3. The $A_{1}$ represents the degree of bending of spin density from the $XY$ plane towards the $Z$-direction (out of plane), whereas $A_{2}$ denotes the intensity of rotation of the $\pi$ orbital within the $XY$ plane. Thus, $A_{1}$ and $A_{2}$ parameters are nearly zero for both the ground and excited states due to minimal participation of $\pi$ orbital and the bending spin density out of the $XY$ plane (Fig. 4(a, b)).
In the relevant ${}^{2}E_{3/2,g/u}$ states split by the external magnetic field, the $A_{1}$ hyperfine interaction mixes $E_{-,\downarrow,g/u}$ with $E_{+,\uparrow,g/u}$. Since $|A_{\|}|\gg|A_{1}|$ and by assuming a positive magnetic field aligned to the symmetry axis of the defect the hyperfine splitting in the lowest energy branch can be approximated as
$A_{g/u}=A_{\|,g/u}+\frac{A_{1,g/u}^{2}}{g_{g/u}\mu_{B}B+A_{\|,g/u}}\text{,}$ (6)
@@
so the actual hyperfine splitting will be weakly magnetic field dependent because of dynamic hyperfine mixing of $E_{+}$ and $E_{-}$ orbitals. As a consequence, the optical transition between hyperfine lines will be also weakly magnetic field dependent: the larger the magnetic field is then the mixing term is more suppressed. By comparing the experimental data and our theory about hyperfine-related splitting in the optical transition we find good agreement. We conclude that the hyperfine couplings are accurately calculated with our method at zero pressure. The hyperfine parameters in the ground and excited states cannot be directly observed in quantum optics measurements so our data can be used for analyzing G4V(--) color centers. We provide further hyperfine data about proximate 13C nuclear spins. These nuclear spins do not reside in the symmetry axis therefore the static and dynamic hyperfine tensors have to be derived for such cases that we present in Appendix B. The results are listed in Table 7. We also studied the quadrupole moment of Ge isotope with $I=9/2$ nuclear spin in GeV(--) for which the dynamical $Q_{1,2}$ may flip the nuclear spin as explained in Refs. Monge et al. (2023); Thiering and Gali (2024). The quadrupole moment ($Q$) can be calculated as
@@
where $\rho$ is the nuclear quadrupole moment, $e$ is the elementary charge, $V_{ZZ}$ is the electric field gradient (EFG) at the nucleus site and $h$ is Planck’s constant Haouas et al. (2016). With using $\rho=-196\times 10^{-31}m^{2}$ (Ref. Pyykkö, 2008) we obtained $Q=-13.4$ MHz, $Q_{1}=-11.5$ MHz and $Q_{2}=-10.3$ MHz.
@@
Figure 6: Illustration of the spin density in $D_{3d}$ symmetry. (a) SiV(--), (b) GeV(--), (c) SnV(--), and (d)PbV(--) defects with zero pressure and (f) SiV(--), (g) GeV(--), (h) SnV(--), and (i) PbV(--) defects under 180 GPa hydrostatic pressure. (e) The scheme of first and second carbon neighbor atoms in black and red spheres, respectively. (j) The spin density of SiV(--) which is representative of all G4V centers in $C_{2h}$ symmetry with no pressure. The electron spin density is depicted with an isosurface value of $5\times 10^{-3}$ $e/$Å.

### v2

@@ III.4 Hyperfine interaction
We computed the hyperfine tensors for the relevant isotopes of the dopant ion and proximate 13C $I=1/2$ nuclear spins that can be optically addressed in a resonant manner [30, 63]. The 13C nuclear spins occur with a 1.1% abundance in natural diamonds. The G4V(--) color centers are often generated by ion implantation, for which the isotope can be well controlled. The non-zero nuclear spin isotopes are 29Si, 73Ge, 117Sn, and 207Pb with respective spins of $I=1/2$, $9/2$, $1/2$, and $1/2$. For 73Ge ($I=9/2$), the nuclear quadrupole moment also contributes to the fine-level structure of the GeV(--) color center, which we also consider in this section.
@@
If these nuclear spins can be well controlled, then they could be applied as quantum memory of the system [64, 23] but may also contribute to extra spin flip-flop processes by reducing the coherence time of the electron spin of G4V(--) centers. Therefore, understanding the full hyperfine tensor and its effects on the system is important in quantum technology applications. As phonons dynamically distort the electronic states via the Jahn-Teller effect, the hyperfine and quadrupole matrix elements become modulated by the vibrational motion, which plays a vital role in spin relaxation processes [65, 66].
@@
We note that the orbital angular momentum is significantly quenched but cannot be entirely neglected [26]. As a consequence, the remaining orbital angular momentum may contribute to the effective dipole-dipole hyperfine coupling. On the other hand, we assume that this contribution is small (e.g., a few percent), and we do not consider it further in this context.
The strong electron-phonon coupling is responsible for relatively short coherence times of the electron spin in G4V(--) defects, even at cryogenic temperatures [67]. Therefore, no electron spin resonance measurements have been reported for G4V(--) defects. The hyperfine structure can be probed optically. Resonant excitation is often performed at the $|{}^{2}E_{3/2,u}\rangle$ level, resulting in $|{{}^{2}\!E_{3/2,u}}\rangle\rightarrow|{{}^{2}\!E_{3/2,g}}\rangle$ optical transitions that appear in the zero-phonon line (ZPL) fluorescence. Because the nuclear spin state is conserved during such optical transitions, only transitions between states with the same nuclear spin projection $m_{I}$ are allowed. This gives rise to transitions such as $|{{}^{2}\!E_{u}^{1}}\rangle\leftrightarrow|{{}^{2}\!E_{g}^{1}}\rangle$ and $|{{}^{2}\!E_{u}^{2}}\rangle\leftrightarrow|{{}^{2}\!E_{g}^{2}}\rangle$. [see Fig. 1(b)]. In the case of $I=1/2$, two hyperfine-resolved lines appear in the spectrum, separated by $A_{\text{PLE}}^{\text{exp}}=A_{\|,u}-A_{\|,g}$.
In typical experimental setups, relatively weak magnetic fields in the range of 1.5–40 G are applied to lift the degeneracy of the Kramers doublets that serve as qubit states. When the magnetic field is aligned along the symmetry axis of the defect, the $|E^{i}_{\{g,u\}\pm}\rangle$ levels are split to first order in perturbation theory by $g\mu_{B}B$ [see Eqs. (C)–(28)], where $g\approx 2$ is the effective $g$-factor in both the ground and excited state manifolds, $\mu_{B}$ is the Bohr magneton, and $B$ is the applied magnetic field strength. Nuclear Zeeman shifts are typically negligible in this regime due to their much smaller magnitude.
@@
Figure 4: The electron spin density of the SiV(--) center with low $C_{2h}$ symmetry in the electronic (a) ground and (b) excited states depicted with the isosurface value of $5\times 10^{-3}$ $e/$Å.
@@
Table 3: Hyperfine parameters of the G4V(--) centers calculated using the HSE06 functional in the $C_{2h}$ symmetry configuration. The parameters $A_{\|}$, $A_{\perp}$, $A_{1}$, and $A_{2}$ are defined in Eqs. (2) and (4), and reported here as reduced values using the reduction factors listed in Table 12. The calculated $A_{\text{PLE}}$ values correspond to the zero-field splitting observed in optical transitions between the $E_{3/2,u}$ and $E_{3/2,g}$ states (see main text for details). Experimental values $A_{\text{PLE}}^{\text{exp}}$ are taken from different studies conducted at various magnetic field strengths. Our calculations indicate that fields in the range of 10–50 G affect $A_{\text{PLE}}$ by less than 0.15%.
@@
29Si $1/2$ $-0.03$ $-0.02$ $2.4$ $5.1$
73Ge $9/2$ $0.22$ $0.52$ $2.3$ 4.6 - -
117Sn $1/2$ $0.1$ $-0.43$ $15.0$ $32.3$
207Pb $1/2$ $0.06$ $0.77$ $-9.7$ $-20.76$
Table 3 lists the calculated hyperfine parameters for dopants in G4V(--) centers. These values were obtained using the HSE06 functional to ensure accurate spin density distributions near the nuclear spins. In all investigated G4V(--) color centers, we find that the spin density around the dopant ion is lower in the electronic excited state compared to the electronic ground state, as illustrated in Figs. 4(a) and (b). This difference can be understood from symmetry considerations: the excited state wavefunction has odd (ungerade) parity, which excludes contributions from $s$-type orbitals since they do not change signs under inversion. In contrast, the ground state wavefunction has even (gerade) parity, allowing $s$-orbital contributions that enhance the Fermi-contact interaction. As a result, the static hyperfine parameters of dopant ions located at the inversion center are more significant in the ground state than in the excited state. Consequently, for dopants with non-zero nuclear spin ($I>0$), the hyperfine splitting $A_{\text{PLE}}$ becomes finite, enabling the optical resolution of hyperfine lines.
As shown in Table 3, the dynamic hyperfine parameters are small but non-zero in both the $|{}^{2}E_{u}\rangle$ and $|{}^{2}E_{g}\rangle$ states. The parameter $A_{1}$ reflects the degree to which the spin density bends out of the $xy$ plane toward the $z$-axis, while $A_{2}$ quantifies the rotational intensity of the $\pi$ orbital within the $xy$ plane. Both $A_{1}$ and $A_{2}$ remain small in the ground and excited states, indicating limited contribution from $\pi$ orbitals and minimal out-of-plane spin density distortion [see Figs. 4(a,b)]. Nevertheless, these components of the hyperfine tensor play a role in shaping the fine structure of the spin levels in G4V(--) centers, as discussed below.
In the relevant $|{}^{2}E_{3/2,\{g,u\}}\rangle$ states split by the external magnetic field and hyperfine interaction terms $A_{\perp,\{g,u\}}$ and $A_{1,\{g,u\}}$ see Eqs. (C) and (C) in the Appendix for details. However, in optical measurements, only the hyperfine splitting ($A_{\text{PLE}}$) of the photoluminescence excitation (PLE) peaks is measured that we show below
$A_{\text{PLE}}(B)=\overset{A_{\text{PLE}}}{\overbrace{A_{\parallel,g}-A_{\parallel,u}}}\pm\left(\frac{A_{\perp,g}^{2}}{\lambda_{g}^{2}}-\frac{A_{\perp,u}^{2}}{\lambda_{u}^{2}}\right)\frac{\mu_{B}g_{S}B}{4}\text{,}$ (5)
where the magnetic field aligned to the symmetry axis. We note that the magnetic field independent terms cancel out or just being negligible see Eq. (29) for details. As a consequence, the optical transition between hyperfine lines will be weakly magnetic field dependent: there will be an extra splitting in the hyperfine peaks at large fields which is very small ($\sim 0.02$ MHz) at $B\sim 10\dots 50$ Gauss that were typically applied in experimental studies [35, 68].
@@
Nevertheless, by comparing the experimental data and our theory about hyperfine-related splitting in the optical transition we find good agreement. We conclude that the hyperfine couplings are accurately calculated with our method at zero pressure. The hyperfine parameters in the ground and excited states cannot be directly observed in quantum optics measurements so our data can be used for analyzing G4V(--) color centers. We provide further hyperfine data about proximate 13C nuclear spins. These nuclear spins do not reside in the symmetry axis therefore the static and dynamic hyperfine tensors have to be derived for such cases that we present in Appendix D. The results are listed in Table 12.
We also studied the quadrupole moment of Ge isotope with $I=9/2$ nuclear spin in GeV(--) for which the dynamical $Q_{1,2}$ may flip the nuclear spin as explained in Refs. [65, 66]. The quadrupole moment ($Q$) can be calculated as
@@
where $\rho$ is the nuclear quadrupole moment, $e$ is the elementary charge, $V_{zz}$ is the electric field gradient (EFG) at the nucleus site and $h$ is Planck’s constant [69]. With using $\rho=-196\times 10^{-31}\mathrm{m}^{2}$ (Ref. 70) we obtained $Q=-13.5$ MHz, $Q_{1}=-5.6$ MHz and $Q_{2}=-5.0$ MHz.
Figure 5: The magnetic field dependency of the ZPL shift is plotted from 0 to 10 T where the magnetic field is aligned with the symmetry axis of the defect. We considered all four possible electronic transitions from excited to the ground states [2] which resulted in four distinct branches of the ZPL for each color center. We performed the calculations for three different hydrostatic pressures of (a) 0, (b) 32, and (c) 180 GPa. Nevertheless, it is expected that the PbV(--) may not exhibit photostable transitions at 180 GPa. The experimental data are depicted in red dots, while the simulated ZPL shifts are represented by blue solid lines.
@@
Figure 7: Illustration of the spin density in $D_{3d}$ symmetry. (a) SiV(--), (b) GeV(--), (c) SnV(--) and (d) PbV(--) defects with zero pressure, and (f) SiV(--), (g) GeV(--), (h) SnV(--) and (i) PbV(--) defects under 180 GPa hydrostatic pressure as obtained by HSE06 functional. (e) The positions of the first and second carbon neighbor atoms are depicted in black and red spheres, respectively. (j) The spin density of SiV(--) representative of all G4V centers in $C_{2h}$ symmetry with no pressure. The electron spin density is depicted with an isosurface value of $5\times 10^{-3}$ $e/$Å.
III.5 Zeeman shift of the optical emission
Next, we investigated the effect of the constant external magnetic field on the optical emission of the G4V centers when the magnetic field is perfectly aligned with the symmetry axis of the defect. The most general microscopic theory of the spin Hamiltonian is given in Refs. [26, 71] that we do not wish to reiterate here. In this microscopic theory, the magnetic field is in part coupled to the orbital moment of the electron that is quenched by the Ham reduction factor (already defined above) and the Stevens orbital reduction factor [72] where the second one is associated with the shape of the orbital (see Ref. 26). We found earlier that the Stevens orbital reduction factor is smaller in the electronic ground state than that in the electronic excited state. We assume that the Steven reduction parameters of the respective electronic states at zero pressure as listed in Appendix D of Ref. 26 are unaltered under hydrostatic pressures. In the simulation of Zeeman shift we go beyond the above applied theory and solve the spin-orbit and electron-phonon Hamitonians simultaneously with the polaronic basis functions $\tilde{\Psi}_{\Gamma}$ with the spin degrees of freedom as
$\left|\tilde{\Psi}_{\Gamma}\right\rangle=\sum_{n,m}\left[c_{nm}^{\chi}\left|e_{g\pm}\right\rangle\left|n,m\right\rangle\left|\chi\right\rangle+d_{nm}^{\chi}\left|e_{g\mp}\right\rangle\left|n,m\right\rangle\left|\chi\right\rangle\right]\text{,}$ (7)
where $\chi$ can be either $\uparrow$ or $\downarrow$ spin state. This solution represents a coupling between spins and phonons and goes beyond the perturbation theory of spin-orbit coupling acting on the polaronic wavefunctions. Here, the subindex $\Gamma$ refers to the total angular momentum of the wavefunction which is either $\frac{3}{2}$ or $\frac{1}{2}$. The parameters of the electron-phonon coupling parameters via JT-effect and the electronic spin-orbit parameter are listed in Tables 2 and 9.
In practical point of view, the orbital moment corrections contribute to the effective $\mathbf{g}$-tensor of the system that may be simply written for the ground state ($g$) and excited state ($u$) as
$\hat{H}^{\text{eff}}_{g,u}=\lambda_{g,u}+{\mu}_{B}\mathbf{S}\mathbf{g}_{g,u}\mathbf{B}\text{,}$ (8)
where $\mu_{B}$ is the Bohr magneton of the electron, $\mathbf{B}$ represents the external homogeneous magnetic field. In this simplified formula, $\lambda_{g,u}$ contains all the electron-phonon related corrections whereas $\mathbf{g}$ amalgamates all the electron-phonon related corrections and orbital moment interaction. We note that the corrections in the Zeeman term specifically occur along the symmetry axis of the defect [26, 71] that causes an observable deviation from the Bohr magneton of the free electron at 2.0023 in the Zeeman shift when the magnetic field is aligned along the symmetry axis of the defect.
With the caveat about the approximations in the microscopic theory of spin Hamiltonian, we calculated the splitting of the ZPL under magnetic fields ranging from 0 to 10 T for G4V centers under hydrostatic pressures of 0, 32, and 180 GPa where the magnetic field is perfectly aligned with the symmetry axis of the defects. We note that it is expected that the PbV(--) may not exhibit photostable optical transitions over 32 GPa but we still calculated it for the sake of complete dataset. Initially, for SiV(--) and SnV(--) centers at zero pressure, we achieved good agreement with experimental data [2, 73] which motivated us to proceed with the calculations under hydrostatic pressures. The results are plotted in Fig. 5. We considered all four possible electronic transitions from excited $E_{u}$ to the ground $E_{g}$ states [2] which resulted in four distinct branches of the ZPL for each color center. The effective g-factors can be read out from the steepness of the curves. We find that the character of the Zeeman splitting under high magnetic fields is basically not altered under high pressure. This is very promising to apply these quantum sensors under extreme conditions.
III.6 Estimation of the electron spin coherence times
In this section, we present our estimation of the electron spin coherence times, a pivotal parameter in quantum information processing applications. We employ the phenomenological theory developed in Ref. 74, which describes the relaxation between $|E_{g\frac{3}{2}}\rangle$ and $|E_{g\frac{1}{2}}\rangle$ orbital states mediated by symmetry breaking $E$ phonons. This model incorporates the electron-phonon coupling strength, $\chi$, as well as the phonon density of states, $\rho$, to describe the underlying relaxation process. The transition between two Kramers states occur via orbital relaxation processes, characterized by the rates $\gamma_{+}$ and $\gamma_{-}$ [see Fig. 1(c)]. In our theoretical framework, we incorporate the ab initio effective spin-orbit splitting which is essential for the estimation of the coherence times. It is important to note that although symmetry-breaking $E$ phonons are mostly responsible for the decoherence processes, their energies are much lower than those quasilocalized $E$ phonons that cause Jahn–Teller-effect. As a consequence, these two regimes are not intertwined; thus, it is legitimate to apply the calculated effective $\lambda$ in this process and independently take the electron-phonon coupling with phonons associated with the decoherence process. Since long-wavelengths phonons responsible for decoherence cannot be directly calculated using our supercell formalism, we extract the parameters $\chi$ and $\rho$ from experimental data that are available at zero pressure. Furthermore we also assume that $\chi$ and $\rho$ do not significantly change in the range of applied hydrostatic pressure. We believe that this simple approximation suffices to sketch the basic trends for the coherence times of the G4V(--) centers.
The temperature as an essential parameter of spin coherence times delineates two regions as found in Ref. 74: (i) the high-temperature region where $T\gg\frac{\hbar\lambda_{g}}{k_{\text{B}}}$ and (ii) the low-temperature region where $T\ll\frac{\hbar\lambda_{g}}{k_{\text{B}}}$. Here, $\hbar$ and $k_{\text{B}}$ are the reduced Planck and Boltzmann constants, respectively, and $\hbar\lambda_{g}$ is the effective spin-orbit strength in energy unit. At high temperatures, the orbital relaxation rate scales as 74
$\gamma_{+}\approx\gamma_{-}\approx\frac{2\pi}{\hbar}\chi\rho\lambda_{g}^{2}k_{\text{B}}T\text{,}$ (9)
where $T$ denotes the temperature. In contrast, at low temperatures, the orbital relaxation rate is given by
$\gamma_{+}$ $\approx 2\pi\chi\rho\lambda_{g}^{3}\exp\left(-\frac{\hbar\lambda_{g}}{k_{\text{B}}T}\right),$ (10)
$\gamma_{-}$ $\approx 2\pi\chi\rho\lambda_{g}^{3}\text{.}$
In this region, acoustic phonons contributing to $\gamma_{+}$ are exponentially suppressed, while the nonradiative relaxation process described by $\gamma_{-}$ persists even at $T=0$ K due to spontaneous decay (e.g., see Ref. 75).
In above equations, $\chi$ and $\rho$ always appear together as a product. Consequently, it is unnecessary to extract these non-calculable terms separately; instead the combination $\chi\rho$ can be considered as a single parameter. We applied the following strategy to extract this parameter from experimental data. We identify the temperature region in terms of the decoherence process at the given temperature and then we apply the respective Eq. (9) for temperature region (i) or Eq. (10) for temperature region (ii). Finally, taking the experimental coherence time $\tau=\gamma^{-1}$ of the G4V(--) center the $\chi\rho$ parameter can be extracted. In this procedure, careful choice of the experimental data is critical as the electron spin coherence time may be limited by proximate nuclear or electron spins in the system at a given temperature. However, our theory only contains the spin-phonon related decoherence, thus the experimental data should be selected from high quality samples to use spin-phonon relaxation related coherence times of G4V(--) centers.
Finally, we note that we apply the calculated effective spin-orbit parameter as obtained by SCAN functional which generally overestimate the experimental data at zero pressure (see Table 2). As a consequence, the estimated coherence times are typically shorter than the experimentally observed values, reflecting the fact that larger spin-orbit splitting leads to reduced coherence times. Nevertheless, we demonstrate below that the estimated values are not far from the experimental ones, thereby, enabling a consistent characterization of the hydrostatic pressure trends using the SCAN functional dataset.
Most of the experimental data is available for SiV(--) center at zero presure. The $\chi\rho$ value increases by raising the temperature from $2.204\times 10^{-29}$ $T=0.1$ K [76] to $1.164\times 10^{-28}$ at $T=3.6$ K [77]. In particular, the measured spin coherence time of SiV(--) with splitting of $\lambda_{\text{exp}}\approx 50$ GHz is reported by values of $\tau^{\text{exp}}_{3.6\text{K}}\approx$115 ns [77] and $\tau^{\text{exp}}_{0.1\text{K}}\approx$13 ms [76]. Our SCAN functional yields $\lambda=70$ GHz which results in $\tau_{3.6\text{K}}\approx$59 ns and $\tau_{0.1\text{K}}\approx$43 ms, respectively, that are in the same order of magnitudes of the experimental data. We further note that two measurement temperatures will fall to different temperature regions in terms of the decoherence process: $T=3.6$ K belongs to region (i) whereas $T=0.1$ K belongs to region (ii). As a consequence, the pressure dependence of the estimated coherence times behaves differently at the two temperatures: $\tau_{3.6\text{K}}$ values decrease with increasing the hydrostatic pressures whereas $\tau_{0.1\text{K}}$ values increase with increasing hydrostatic pressures (Table 4).
Spin coherence time of 20 ms was reported [78] for GeV(--) with $\lambda_{\text{exp}}=181$ GHz at $T=300~{}\text{mK}$. The extracted $\chi\rho$ is $9.402\times 10^{-31}$ and it belongs to the temperature region (ii). In this case, the estimated spin coherence time increases from 49 ms (zero pressure) up to 2.63 s at hydrostatic pressure of 180 GPa (Table 5).
Next, we study the coherence times of SnV(--) based on the available experimental data [13] at temperatures of 2.9 K and 6 K where $\lambda_{\text{exp}}=850$ GHz and the decoherence process goes in the region (ii) at both temperatures. Subsequently, $\chi\rho$ is $1.175\times 10^{-29}$ at 2.9 K and $1.986\times 10^{-29}$ at 6 K. In our simulations, $\tau_{2.9\text{K}}$ values increase from 545 ns to 1.3 $\mu$s whereas $\tau_{6\text{K}}$ values decrease from 55 ns to 35 ns (Table 6).
To our knowledge, no coherence times for PbV(--) has been reported. However, it is predicted to exhibit a millisecond spin coherence time at 9 K [18] with $\lambda_{\text{exp}}=3914$ GHz which is located in temperature region (ii). As a result, $\chi\rho=3\times 10^{-34}$ may be extracted. Finally, the estimated spin coherence times raised from 1.29 ms to 1.85 ms at 9 K from zero to 32 GPa hydrostatic pressures at photostable region (Table 7).
Table 4: Hydrostatic pressure (P) dependency of spin coherence time ($\tau$) at the given temperatures at zero to 180 GPa pressure for SiV(--) center in diamond as obtained by SCAN functional. These parameters with $T=0.1$ K results in ($T\ll\frac{\hbar\lambda_{g}}{k_{\text{B}}}$) so Eq. 10 is applied whereas it yields ($T\gg\frac{\hbar\lambda_{g}}{k_{\text{B}}}$) with $T=3.6~{}\text{K}$ so Eq. 9 is applied.
P (GPa) $\lambda$ (GHz) $\tau$ ($\tau_{\text{exp}}$) (ms) $\tau$($\tau_{\text{exp}}$) (ns)
$T=0.1K$ $T=3.6K$
0 70 43 (13111Ref. 77) 59 (115222Ref. 76)
32 76 64 50
72 82 98 43
120 88 147 38
180 93 220 34
Table 5: Hydrostatic pressure (P) dependency of spin coherence time ($\tau$) at $T=0.3$ K at zero to 180 GPa pressure for GeV(--) center in diamond as obtained by SCAN functional. These parameters results in ($T\ll\frac{\hbar\lambda_{g}}{k_{\text{B}}}$) so Eq. 10 is applied.
P (GPa) $\lambda$ (GHz) $\tau$ ($\tau_{\text{exp}}$) (ms)
0 223 49 (20111Ref. 78)
32 256 106
72 294 278
120 332 762
180 376 2626
Table 6: Hydrostatic pressure (P) dependency of spin coherence time ($\tau$) at the given temperatures at zero to 180 GPa pressure for SnV(--) center in diamond as obtained by SCAN functional. These parameters with $T=2.9$ K results in ($T\ll\frac{\hbar\lambda_{g}}{k_{\text{B}}}$) so Eq. 10 is applied whereas it yields ($T\gg\frac{\hbar\lambda_{g}}{k_{\text{B}}}$) with $T=6~{}\text{K}$ so Eq. 9 is applied.
P (GPa) $\lambda$ (GHz) $\tau$ ($\tau_{\text{exp}}$) (ns) $\tau$ ($\tau_{\text{exp}}$) (ns)
$T=2.9~{}\text{K}$ $T=6~{}\text{K}$
0 915 545 (540111Ref. 18) 55 (59111Ref. 18)
32 1068 608 45
72 1224 724 40
120 1407 946 37
180 1586 1291 35
55footnotetext: Ref. 13
Table 7: Hydrostatic pressure (P) dependency of spin coherence time ($\tau$) at $T=9~{}K$ at zero to 180 GPa pressure for PbV(--) center in diamond as obtained by SCAN functional. These parameters results in ($T\ll\frac{\hbar\lambda_{g}}{k_{\text{B}}}$) so Eq. 10 is applied.
P (GPa) $\lambda$ (GHz) $\tau$ ($\tau_{\text{exp}}$) (ms)
0 4436 1.29 (1.0111Ref. 18)
32 5065 1.85
72 5744 2.87
120 6387 4.54
180 6593 5.29